OpenStudy (anonymous):

the length of a rectangle is 3 ft more than twice the width. the area is 90 square ft. what are the dimnesions of the rectangle.

OpenStudy (anonymous):

width = x ; length = 2x+3ft x(2x+3)=90 2x^2 + 3x = 90 using quadratic formula discriminant=729 x=(-3+27)/4=6ft (taking 27 +ve coz dimensions cannot be -ve) so width = 6ft and lenght 15ft